# KunServe: Efficient Parameter-centric Memory Management for LLM Serving

### v4

@@ Introduction
A key takeaway from Figure 5 is that the performance loss is highly correlated with the number of instances involved in processing a request, i.e., pipeline stages. Thus, we design a greedy-based parameter dropping algorithm by grouping as few instances as possible for meeting the above requirement (3), i.e., minimize performance loss.
Algorithm 6 shows the details of our method. Upon throttling, we first compute the memory demand of all queuing requests ($R$). Afterward, we group instances belonging to a pipelined execution ($G$), and then iteratively drop parameters to free more space (line 3). Note that during the initialization time, each pipelined group only has one instance. To minimize the number of instances within each pipelined execution group, we first drop parameters at groups with fewer instances (line 2, the priority is reversed to the group size). For the selected groups, they will drop half of the redundant parameters each and form a new pipelined group (line 5–8), and the freed memory is updated (line 9). The iteration continues until the memory requirement is satisfied or we fail to find a drop plan (line 3). The complexity of the plan generation is $O(N\log{N})$, so we can quickly execute it online even with a large number of instances, thus satisfying (1).
@@
Figure 7: The pagedattention GPU kernel signature [2] (a) and (b) CUDA virtual memory management APIs [4].
@@
Local instance memory management. A key challenge of executing the drop plan at each instance is how to allow existing attention kernels to use the freed parameter memory. As shown in Figure 7 (a), the kernels are written with a single static memory layout, e.g., [kcacheaddr, kcacheaddr + numblocks * blocksize], not multiple virtual memory ranges provisioned dynamically. One possible solution is to rewrite these kernels to suit the new memory layout. However, efficiently rewriting LLM kernels is non-trivial due to the complex and evolving nature of LLM kernels. Simple rewrites lead to performance drops that require months of iterative development to optimize [36].
@@
Due to the auto-regressive nature of LLMs, after the parameter drop, we cannot simply execute ongoing decode requests because the execution needs a one-to-one mapping between the parameters and the KVCache. For example, suppose a request has executed on instance A, and A forms a group with instance B due to memory throttling. After the drop, A will only have half of the parameters, so the computation based on KVCache generated by the dropped parameters cannot proceed unless we recompute these KVCache. However, the recomputation is expensive because a request may have entered the decode phase for a while, so many KVCaches and tokens need to be recomputed.
Network-based KVCache exchange. We choose to exchange the KVCache through the network to avoid recomputation. The KVCache is exchanged because after A and B have formed a group, ongoing requests on A need to transfer their KVCache to B, while B needs to do so vice versa. During the exchange, these requests will be stalled. The stall time is manageable because the network between instances like RDMA is sufficient for transferring the KVCache quickly. For example, KVCache exchange typically introduces 1–2 s stall time on our 200 Gbps network. This means a 10 ms increase at most in TPOT metric of a response with 200 decode tokens.
Note that during the stall, we can still schedule new requests queued due to memory throttling to fully utilize the GPUs. While in principle we can leverage techniques like attention offloading (also called model-attention disaggregation) [14] to concurrently execute stalled requests during the KVCache exchange, we found the excessive complexity of the implementation is not worth the effort.
Coordinated KVCache exchange. Although straightforward, KVCache exchange could block new request execution if not designed carefully. This is because the exchange competes for bandwidth with activation transfers in pipelined execution. Since the exchange time is much longer than forwarding the activation, When the activation is waiting for the exchange to finish, it will leave the GPUs idle, causing non-negligible performance loss. Observing that the activation transfer is much smaller yet more critical, we design a coordinated KVCache exchange mechanism to prioritize the activation transfer. Specifically, we will transfer KVCache in finer-grained chunks such that the transfer of a time is similar to the execution time of a pipeline stage. After transferring one chunk, we will check whether there would be activation transfer. If so, we will first transfer the activation and then resume the KVCache transfer.
@@
Figure 8: An illustration of pipeline execution bubbles caused by imbalanced execution time of microbatches.
@@
Figure 9: (a) An illustration of serving requests to execute. (b) The imbalanced batch execution time of existing chunking method. (c) A balanced formulated batch configuration.
A preliminary on the state-of-the-art pipeline microbatch formulation. Modern pipeline implementations rely on chunked prefill to reduce pipeline bubbles. Specifically, they [6, 27] form microbatches in a token-count-based manner, which balances the execution time of different microbatches by ensuring each microbatch has a similar number of tokens. As shown in Figure 17 (a), suppose 5 requests (R0–R4) arrive at an instance in turn, and the budget for each microbatch is 4 tokens. The scheduler first merges incoming requests into one microbatch (R0–R3 in (b)). R4 itself forms another microbatch (B1). Note that if R4 exceeds the budget, the scheduler will chunk it into two segments for execution.
Inefficiency of token-count-based chunking. A key problem is that the microbatch execution time is not linearly proportional to the total token count, because the attention computation of each request is quadratic to its token count, as shown in Figure 17 (b). Moreover, if a request is chunked into two parts, the latter chunk is slower than the former even when the tokens are the same, because the latter chunk has to additionally compute the attention with the former chunk.
Figure 10: An illustration of how lookahead batch formulation recursively generate balanced microbatches.
The lookahead batch formulation. Fortunately, under bursts, we have sufficient requests queued. Thus, we can re-form the microbatches across them by looking ahead all requests queued. However, finding the optimal balance plan is NP-hard (detailed in §A due to space limitation). Thus, we propose a heuristic divide-and-conquer algorithm to quickly find a balanced batch configuration.
Figure 11: The pseudocode of the divide-and-conquer microbatch formulation algorithm.
The method works as a two-step process. First, we adopted a retrofitted cost model to precisely estimate the execution time of a microbatch. Second, we recursively generate the microbatch configurations according to the cost model. Specifically, balancing can be done by looking ahead all tokens to be chunked in a recursive manner, as shown in Figure 10. The initial batch contains a single microbatch with all tokens, which is then recursively split into two cost-balanced microbatches until it reaches balanced setup.
@@
A key to the effectiveness the above algorithm is how to accurately estimate the execution time (i.e., cost) of a microbatch. We acquire the cost model using a bottom-up approach: we first model the cost of executing a chunk of a request, then we sum the cost of all chunks in a microbatch as its cost. Specifically, suppose we have a microbatch set $\mathcal{B}$, denoted by =$\mathcal{B}=\{b_{\_}1,b_{\_}2,\ldots,b_{\_}m\}$, The chunks are chunked from a request set of size $n$, denoted by =$\mathcal{R}=\{r_{\_}1,r_{\_}2,\ldots,r_{\_}n\}$. The cost of a chunk $c_{\_}{ij}$$\text{cost}_{\_}{c_{\_}{ij}}$ can be formulated as follows:
$\text{cost}_{\_}{c_{\_}{ij}}=\alpha\,\left(\overbrace{p_{\_}{ij}c_{\_}{ij}}^{\text{prefix-attn}}+\overbrace{\frac{c_{\_}{ij}^{2}+c_{\_}{ij}}{2}}^{\text{self-attn}}\right)+\beta\,\overbrace{c_{\_}{ij}}^{\text{mlp}}+\overbrace{\gamma}^{\text{other}}$ (1)
The equation consists of three parts: the cost to compute attention with previous tokens (prefix-attn); the cost to compute attention with the chunk itself (self-attn); the cost to forward MLP layers for tokens; and other costs that are amortized by a batch. The prefix tokens of each chunk can be calculated as $p_{\_}{ij}=\sum_{\_}{k=1}^{j-1}c_{\_}{ik}$. The prefix-attn and self-attn terms model the quadratic cost of attention computation missed by existing cost models. The model depends on several hyperparameters that can be profiled offline, similar to existing works like DistServe [50] and NanoFlow [51]. While we both leverage profiling for performance modeling, NanoFlow does not consider the self-attn while DistServe does not take prefix-attn into account.
Given the cost of each chunk, we can sum all the cost of chunks in a microbatch to get the cost of the microbatch:
$b_{\_}k=\left\{c_{\_}{ij}\mid x_{\_}{ij}=k\land c_{\_}{ij}>0\right\},\quad\forall k\in\{1,\ldots,m\}$ (2)
$\text{cost}_{\_}{b_{\_}k}=\sum_{\_}{\begin{subarray}{c}(i,j)\\ x_{\_}{ij}=k\end{subarray}}\text{cost}_{\_}{c_{\_}{ij}}$ (3)
@@
Empirically, our cost model models the execution time of microbatch, efficiently. As a result, the pipelined executed with our lookahead formulation can significantly reduce the execution bubbles (see Figure 14).
@@
Discussion: the generality of lookahead batch formulation. While in principle, we could also apply lookahead batch formulation to general LLM serving with pipeline execution, it has one obstacle that the formulation assumes a sufficient number of requests queued to “lookahead” to be effective. Under normal serving without bursts, waiting for requests to be looked ahead may add additional latency, which we leave possible solutions as a future work.
Dynamic restore, scaling and fault tolerance
@@
Dynamic parameter restoration. While dynamic parameter drop described in §4.1 can free up memory for new requests under memory throttling, the pipelined execution is not optimal under normal execution because (1) pipelined execution suffers from more frequent weight loading and (2) it has bubbles. Normal execution cannot simply apply our lookahead scheduling described in §4.3 because there are insufficient numbers of requests to balance.
@@
To this end, KunServe dynamically restores parameters to return to a normal non-pipelined execution once the throttling fades away. Specifically, when the monitor detects that the total KVCache usage is below a threshold, it triggers a restoration process by loading the dropped parameters back to the GPUs. Currently we use a simple threshold where the memory usage is below 50 % of the GPU (without drop). The missing parameters are pulled from instances whenever possible using the network between instances.
@@
Autoscaling. A limitation of parameter centric memory management is that the memory freed for KVCache is limited, so KunServe cannot handle throttling caused by continuous request bursts. Fortunately, such bursts are rare in practice. More importantly KunServe can sustain 1.5 $\times$ longer than current systems without affecting SLOs even under extreme unrealistic bursts (§5.5).
To cope with this limitation, we autoscale the systems by allocating new GPUs and deploying new instances on them, similar to prior work [41, 20], when we detect a continuous increase in the memory usage after the drop. The extra time gained by our parameter drop (33s) is sufficient to hide the relatively slow autoscaling process [35].
@@
Testbed. We evaluate KunServe on two clusters listed in Table 2. Cluster A has one GPU per server so it is typically used for running small models (e.g., 14 B models). On the other hand, cluster B has multiple GPUs per server interconnected with fast NVLink, so it is suitable for running larger models (e.g., 72 B models) with tensor parallelism.
@@
Evaluated models. Similar to prior works [6, 35, 50], we choose open-source models with leading accuracy: Qwen-2.5-14B and Qwen-2.5-72B [42]. Both models adopt GQA [9] to reduce KVCache size while maintaining high accuracy. We don’t choose models with huge KVCache assumption (e.g., models with MHA [43]) that could easily throttle memory to show KunServe’s effectiveness in the state-of-the-art models. Table 1 lists instance configurations of each model. For the 72B model, we use tensor parallelism to serve requests on multiple GPUs.
@@
Evaluated traces and datasets. Since memory throttling is sensitive to the request arrival pattern, we use a real-world trace BurstGPT [45] with known request arrival information (i.e., the invocation time of each request) as our main evaluated application. Following the guide of BurstGPT, we scale BurstGPT’s RPS to fit the serving capacity of our testbed using a scaling method that preserves the temporal pattern of the trace. Specifically, we upscale the trace with TraceUpscaler [38], and ensure that the average memory demand is lower than 60% of the total memory during the entire evaluation of the trace.
@@
Besides the arrival pattern, LLM serving is also sensitive to the input and output length of requests. Thus, given the trace, we further evaluate requests from representative datasets representing different scenarios, similar to prior works [31, 29]:
@@
  • BurstGPT. It is the original dataset of BurstGPT [45], representing a conversion workload so both TTFT and TPOT are important. The average input and output lengths are 642 and 262, respectively.
@@
  • ShareGPT. ShareGPT [3] is another popular chatbot dataset that is widely evaluated on [41, 50, 47, 6]. Its input and output lengths are longer than BurstGPT, representing a workload that is more sensitive to GPU memory provisioning. The maximal input length is 4K, and the average input and output lengths are 1,660 and 373, respectively. Like BurstGPT, low TTFT and TPOT are both important for benchmark using this dataset.
@@
  • LongBench. LongBench [13] is another popular dataset used for evaluating document summarization tasks [50], e.g., summarizing news, articles and scientific papers. The average input length is 5.9 K and the average output length is 499. Since the user expects a quick response to the summarized content, TTFT is also important.
@@
Figure 12: First column: the memory usage pattern of KunServe. Second column: the mean TTFT during the evaluation. Third column: the throughput during the evaluation.
@@
Table 2: Testbed. $s$ is the number of servers and $g$ is the number of GPUs per host. The scaleup and scaleout here means scale-up network and scale-out network, respectively. The reported bandwidth is unidirectional.
Baselines. We compared with the state-of-the-art LLM serving systems with various techniques to cope with memory throttling. For all systems, we have carefully tuned their configurations to meet the optimal performance without memory throttling. We have also enabled all known serving optimizations to these systems even though the vanilla systems are not optimized (e.g., InferCept [5]). For those with our optimizations, we have calibrated that our optimizations enabled better performance than the original open-sourced codebase. More specifically, our baselines are:
  • vLLM (default + PP) [27]. We compare two configurations of vLLM (release v0.6.3): The default configuration stores the entire parameters on each instance, while pipelined parallelism (PP) further frees half of the parameters on each instance and leverages PP to execute requests across two instances. This setup frees up more memory for KVCache, but it also introduces pipelined execution overhead. By default, vLLM uses recomputation to cope with memory throttling.
  • InferCept [5]. InferCept designs an optimized swap mechanism that eliminate IO idle time atop vLLM. We tried to compare its open-sourced version, but found its performance is 1.2–5.1 $\times$ slower in TTFT and 1.2–1.9 $\times$ in TPOT than the chosen vLLM release even without memory throttling. This is because it was implemented on an old version of vLLM (v0.2.0), where important optimizations (e.g.,  FlashAttention/FlashInfer kernels [15, 18], chunked prefill [6]) are missing. We integrate our scheduler and attention backend into the original InferCept for a fair comparison.
  • Llumnix [41]. Llumnix adopts load balancing to cope with memory throttling of an instance, and migrates KVCache between instances to free sufficient memory in case of insufficient memory even with load balancing. We compared with the latest version of Llumnix (release v0.1.0).
@@
Figure 13: The end-to-end latency results. Column from 1 to 4 is the end-to-end metrics of different workloads. The last column is the SLO violation of TTFT and TPOT with different SLO scales.
@@
The second column of Figure 12 presents how the mean TTFT changes over time given a measured time window (e.g., 100s), and Figure 13 presents the zoomed-in view of the P50 and P99 latencies when evaluating different workloads on different models. First, KunServe has 12.7–72.2 $\times$ faster P99 TTFT than other baselines, because it frees up sufficient memory under memory throttling, which enables requests queued in other systems to be served with a larger batch size. For other systems, they either suffer from recomputation overhead (vLLM), or queuing overhead waiting for swapping (InferCept) or migration (Llumnix) under memory throttling. Specifically, the timeline plotted in Figure 12 clearly shows that the TTFT increase coincides with the increased KVCache demands (the first column in Figure 12).
@@
Compared to other baselines, KunServe trades a little increase in P50 TPOT, and P99 TPOT because it executes requests in a larger batch to eliminate queuing. For example, in LongBench-14B workload, the P50 TPOT of KunServe is 15.8–22.7% higher than other baselines. We believe it is a reasonable trade-off because such increases are still within the SLOs of targeted applications, which we describe next. Interestingly, KunServe even has a little P50 TTFT improvement in the LongBench workload. This is because the long and diverse input of requests in this workload makes the system more prone to memory throttling caused by severe memory fragmentation [41]. Thus, the many queued requests affect normal requests.
SLO attainment. SLO is an important metric for serving systems [29, 50], which defines the maximum acceptable latency for a request. Requests whose latency exceeds the SLO are not useful because users may abandon them [37]. Because different applications have different maximum acceptable latency requirements (SLOs), we evaluate the SLO violation of all systems under different SLO scale factors, similar to previous works [29, 35, 40, 50].
Specifically, the last column of Figure 13 shows the SLO violation of all systems with different SLO scale factors, where a scale factor of $N$ means that the maximal tolerable latency is $N$ times the P50 latency of the best baseline. To help understand how the reduced SLO violations of KunServe benefits end-to-end applications, we also mark the typical scale for our evaluated applications, i.e., we set 5 for chat—a tight SLO as it requires quick responsiveness, while for document summarization, we set a looser factor of 10, following previous works [50]. We can see that KunServe achieves 7.2—12.8% average SLO violation reductions on various workloads, and more importantly, it almost eliminates all violations with a scale larger than 4 for all workloads. Other baselines cannot eliminate SLO violations even with an extremely loose factor of 10 because during bursts, there are considerable numbers of queued requests suffering from 45—840 $\times$ tail latency increases.
Multi-GPU model performance. Due to space limitations, we only present the results of multi-GPU model (Qwen-2.5-72B) on LongBench dataset. Results on other datasets are similar. As shown in Figure 12 and Figure 13, the trend is similar to single-GPU models: KunServe reduces the P99 latency by 8.4–11.9 $\times$ compared to other baselines, at the cost of a slight (18.3–22.7%) increase in P50 TPOT and P99 TPOT. Multi-GPU model has a similar result because each instance (containing multiple GPUs) as a whole can be viewed as a single logical GPU. Multi-GPU model even has a larger benefit from dropping parameters because the relative ratio of parameter memory is large, see Table 1.

### v5

@@ Introduction
A key takeaway from Figure 5 is that the performance loss is strongly correlated with the number of instances involved in processing a request, i.e., pipeline stages. Thus, we design a greedy-based parameter dropping algorithm by grouping as few instances as possible to minimize performance loss.
[…]
Upon overloading, we first compute the memory demand of all queued requests ($R$) and enter line 1. Afterward, we iteratively group instances and then drop parameters to free more space (lines 3–9). For example, if there are three groups with sizes of 1, 2, and 3, we will select the two groups with sizes of 1 and 2 to form a new group (lines 5–6). For the selected groups, we drop a copy of the redundant parameters (line 7) and update the available memory (line 9). At the end of the iteration, the selected two groups are merged into a new group and inserted back into the priority queue (line 8).
The iteration continues until the memory requirement is satisfied or it fails to find a drop plan (line 3). In case we cannot find a plan, we fallback to the KVCache-centric solution to ensure continuous execution and autoscale the instance numbers. The complexity of the plan generation is $O(N\log{N})$, so we can quickly execute it online even with a large number of instances.
@@
Figure 7. (a) The GPU kernel signature of the pagedattention kernel (vllm-code, ). (b) CUDA virtual memory management APIs (cuda-memory, ).
@@
Local instance memory management.    A key challenge of executing the drop plan at each instance is how to allow existing attention kernels to use the freed parameter memory. As shown in Figure 7 (a), the kernels are written with a single static memory layout, e.g., [kcacheaddr, kcacheaddr + numblocks * blocksize], not multiple virtual memory ranges provisioned dynamically. One possible solution is to rewrite these kernels to suit the new memory layout. However, efficiently rewriting LLM kernels is non-trivial due to the complex and evolving nature of LLM kernels. Simple rewrites lead to performance drops that require months of iterative development to optimize (DBLP:journals/corr/abs-2405-04437, ).
@@
Because the KVCache has a one-to-one mapping with the model parameters, we cannot simply execute ongoing decode requests due to the lack of KVCache. For example, suppose a request has executed on instance A, and A has formed a group with instance B due to memory overloading. After the drop, A will only have parameters of layers 0–4, while B will have layers 5–7. Hence, B cannot directly execute the 5–7 layers of a request originally on A because the required KVCache is on A. Similarly, A cannot execute the 0–4 layers of a request originally on B. One intuitive solution is to recompute the KVCache on B. This is expensive since it causes queued requests to wait for the recomputation even without considering the recomputation time.
Network-based KVCache exchange.    We choose to exchange the KVCache through the network to avoid recomputation. The KVCache is exchanged because after A and B have formed a group, ongoing requests on A need to transfer their KVCache to B, while B needs to do the same vice versa. A drawback of the exchange is that the requests with the exchanged KVCache will be stalled during the exchange, which we found to be acceptable in practice. This is because the network between instances such as RDMA is sufficient for transferring the KVCache quickly. For example, KVCache exchange typically introduces 1–2 s stall time on our 200 Gbps network. This means a 10 ms increase at most in the TPOT metric of a response with 200 decode tokens.
Note that during the stall, we can still schedule new requests queued due to memory overloading to fully utilize the GPUs. While in principle we can leverage techniques like attention offloading (also called model-attention disaggregation) (DBLP:journals/corr/abs-2405-01814, ) to concurrently execute stalled requests during the KVCache exchange, we found the excessive complexity of the implementation is not worth the effort.
Coordinated KVCache exchange.   Although straightforward, KVCache exchange could block new request if not implemented properly, because the exchange competes for bandwidth with activation transfers in pipelined execution. Since the exchange time is much longer than forwarding the activation, When the activation is waiting for the exchange to finish, it will leave the GPUs idle, causing non-negligible performance loss. Observing that the activation transfer is much smaller yet more critical, we design a coordinated exchange mechanism to prioritize the activation transfer. Specifically, we transfer KVCache in finer-grained chunks such that the transferring a chunk takes similar time to executing a pipeline stage. After transferring one chunk, we will check whether there will be activation transfer. If so, we pause the KVCache transfer and let the activation transfer go first.
@@
Figure 8. An illustration of pipeline execution bubbles caused by imbalanced execution time of microbatches.
Figure 9. (a) An illustration of serving requests to execute. (b) The imbalanced batch execution time of existing chunking method. (c) A balanced formulated batch configuration.
@@
A preliminary on the state-of-the-art pipeline microbatch formulation.    Modern pipeline implementations rely on chunked prefill to reduce pipeline bubbles. Specifically, they (sarathi, ; vllm, ) form microbatches in a token-count-based manner, which balances the execution time of different microbatches by ensuring each microbatch has a similar number of tokens. As shown in Figure 9 (a), suppose 5 requests (R0–R4) arrive at an instance in turn, and the budget for each microbatch is 4 tokens. The scheduler first merges incoming requests into one microbatch (R0–R3 in (b)). R4 itself forms another microbatch (B1). Note that if R4 exceeds the budget, the scheduler will chunk it into two segments for execution.
Inefficiency of token-count-based chunking.    A key issue is that the microbatch execution time is not linearly proportional to the total token count, because the attention computation of each request is quadratic to its token count, as shown in Figure 9 (b). Moreover, if a request is chunked into two parts, the latter chunk is slower than the former even when the tokens are the same, because the latter chunk has to additionally compute the attention with the former chunk.
Figure 10. An illustration of how lookahead batch formulation recursively generate balanced microbatches.
The lookahead batch formulation.    Fortunately, under bursts, we have sufficient requests queued. Thus, we can re-form the microbatches across them by looking ahead at all requests queued. To efficiently find the balanced microbatch configuration, we propose a heuristic divide-and-conquer algorithm.
Figure 11. The pseudocode of the divide-and-conquer microbatch formulation algorithm.
Our method works in two steps. First, we adopted a retrofitted cost model to precisely estimate the execution time of a microbatch. Second, we recursively generate the microbatch configurations according to the cost model. Specifically, balancing can be done by looking ahead all tokens to be chunked in a recursive manner, as shown in Figure 10. The initial batch contains a single microbatch with all tokens, which is then recursively split into two cost-balanced microbatches until it reaches a balanced setup.
@@
[…]
Our model depends on several hyperparameters (e.g., $\alpha$) that can be determined through offline profiling: before the system is deployed for serving, we run multiple inference samples offline, collect their execution times, and then use the least squares method (least-square, ) to determine all hyperparameters.
Given the cost of each chunk, we can sum all the costs of chunks in a microbatch to get the cost of the microbatch:
[…]
@@
Empirically, our cost model accurately models the execution time of a microbatch for common sequence lengths in Figure 15. As a result, the pipelined execution with our lookahead formulation can significantly reduce the execution bubbles (see Figure 14).
@@
Discussion: the generality of lookahead batch formulation and cost model.    While in principle, we could also apply lookahead batch formulation to general LLM serving with pipeline execution, it has one obstacle that the formulation assumes a sufficient number of requests queued to “lookahead” to be effective. Under normal serving without bursts, waiting for requests to be looked ahead may add additional latency, which we leave possible solutions as a future work.
Besides, readers may findEq. 1 still has a part that has a linear correlation with the number of tokens (FFN), so if the cost is dominated by FFN, existing token-count-based cost models may suffice. We argue that our retrofitted cost model is still important because the quadratic terms (prefix-attn and self-attn) would become significant when the token count increases (e.g., for requests with more than 4K tokens, which are common in real-world workloads (longbench, ), see §5.1), so existing works can leverage our model for a more accurate estimation of microbatch execution time.
Dynamic restore and fault tolerance
@@
Dynamic parameter restoration.  While dynamic parameter drop described in §4.1 can free up memory for new requests under memory overloading, the pipelined execution is not optimal under normal execution because (1) pipelined execution suffers from more frequent weight loading and (2) it has bubbles. Normal execution cannot simply apply our lookahead scheduling described in §4.3 because there are insufficient numbers of requests to balance.
@@
To this end, KunServe dynamically restores parameters to return to a normal non-pipelined execution once the overloading fades away. Specifically, when the monitor detects that the total KVCache usage is below a threshold, it triggers a restoration process by loading the dropped parameters back to the GPUs. Currently we use a simple threshold where the memory usage is below 50 % of the GPU (without drop). The missing parameters are pulled from instances whenever possible using the network between instances.
@@
Testbed.    We evaluate KunServe on two clusters listed in Table 2. Cluster A has one GPU per server so it is typically used for running small models (e.g., 14 B models). Cluster B has multiple GPUs per server interconnected with fast NVLink, so it is suitable for running larger models (e.g., 72 B models) with tensor parallelism.
@@
Evaluated models.    Similar to prior works (sarathi, ; splitwise, ; distserve, ), we choose open-source models with leading accuracy: Qwen-2.5-14B and Qwen-2.5-72B (qwen2.5, ). Both models adopt GQA (gqa, ) to reduce KVCache size while maintaining high accuracy. We do not choose models with huge KVCache usage (e.g., models with MHA (DBLP:conf/nips/VaswaniSPUJGKP17, )) that could easily exhaust GPU memory—though KunServe is more effective when serving such models. This is because these models are being replaced by more KVCache-efficient variants. Table 1 lists instance configurations of each model. For the 72B model, we use tensor parallelism to serve requests on multiple GPUs.
@@
Evaluated traces and datasets.    Since memory overloading is sensitive to the request arrival pattern, we use a real-world trace BurstGPT (burstgpt, ) with known request arrival information (i.e., the invocation time of each request) as our main evaluated application. Following the guide of BurstGPT, we scale BurstGPT’s RPS to fit the serving capacity of our testbed using a scaling method that preserves the temporal pattern of the trace. Specifically, we upscale the trace with TraceUpscaler (traceupscaler, ), and ensure that the average memory demand is lower than 60% of the total memory during the entire evaluation of the trace.
@@
Besides the arrival pattern, LLM serving is also sensitive to the input and output length of requests. Thus, given the trace, we further evaluate requests from representative datasets representing different scenarios, similar to prior works (spotserve, ; DBLP:conf/osdi/LiZZL00HCZGS23, ):
@@
  • BurstGPT. It is the original dataset of BurstGPT (burstgpt, ), representing a conversion workload so both TTFT and TPOT are important. The average input and output lengths are 642 and 262, respectively.
@@
  • ShareGPT. ShareGPT (sharegpt, ) is another popular chatbot dataset that is widely evaluated on (llumnix, ; distserve, ; loongserve, ; sarathi, ). Its input and output lengths are longer than BurstGPT, representing a workload that is more sensitive to GPU memory provisioning. The maximal input length is 4K, and the average input and output lengths are 1,660 and 373, respectively. Like BurstGPT, low TTFT and TPOT are both important for benchmark using this dataset.
@@
  • LongBench. LongBench (longbench, ) is another popular dataset used for evaluating document summarization tasks (distserve, ), e.g., summarizing news, articles and scientific papers. The average input length is 5.9 K and the average output length is 499. Since the user expects a quick response to the summarized content, TTFT is also important.
@@
Figure 12. First column: the memory usage pattern of KunServe. Second column: the mean TTFT during the evaluation. Third column: the throughput during the evaluation.
@@
Table 2. Testbed. $s$ is the number of servers and $g$ is the number of GPUs per host. The scaleup and scaleout here means scale-up network and scale-out network, respectively. The reported bandwidth is unidirectional.
Baselines.    We compared with the state-of-the-art LLM serving systems with various techniques to cope with memory overloading. For all systems, we have carefully tuned their configurations to meet the optimal performance without memory overloading. We have also enabled all known serving optimizations to these systems even though the vanilla systems are not optimized (e.g., InferCept (infercept, )). For those with our optimizations, we have calibrated that our optimizations enabled better performance than the original open-sourced codebase. More specifically, our baselines are:
  • vLLM (default + PP) (vllm, ). We compare two configurations of vLLM (release v0.6.3): The default configuration stores the entire parameters on each instance, while pipelined parallelism (PP) further frees half of the parameters on each instance and leverages PP to execute requests across two instances. This setup frees up more memory for KVCache, but it also introduces pipelined execution overhead. By default, vLLM uses recomputation to cope with memory overloading. We compared the vLLM with swapping to InferCept described below.
    Before the evaluation, we carefully tuned the configurations of vLLM. Specifically, we tuned the block size to achieve the best performance under our setup. We chose 64 because (1) it is small enough to avoid memory fragmentation while (2) it is sufficiently large to achieve good performance (flashinfer, ).
  • InferCept (infercept, ). InferCept designs an optimized swap mechanism that eliminates IO idle time atop vLLM. We tried to compare its original open-sourced version, but found its performance is 1.2–5.1 $\times$ slower in TTFT and 1.2–1.9 $\times$ in TPOT than the chosen vLLM release even without memory overloading. This is because it was implemented on an old version of vLLM (v0.2.0), where important optimizations (e.g.,  FlashAttention/FlashInfer kernels (flashattention2, ; flashinfer, ), chunked prefill (sarathi, )) are missing. Therefore, we integrated our scheduler and attention backend into the original InferCept for a fair comparison.
  • Llumnix (llumnix, ). Llumnix adopts load balancing to cope with memory overloading of an instance, and migrates KVCache between instances to free sufficient memory in case of insufficient memory even with load balancing. We compared with the latest version of Llumnix (release v0.1.0).
@@
Figure 13. The end-to-end latency results. Column from 1 to 4 is the end-to-end metrics of different workloads. The last column is the SLO violation of TTFT and TPOT with different SLO scales.
@@
The second column of Figure 12 presents how the mean TTFT changes over time given a measured time window (e.g., 100s), and Figure 13 presents the zoomed-in view of the P50 and P99 latencies when evaluating different workloads on different models. First, KunServe has 12.7–72.2 $\times$ faster P99 TTFT than other baselines, because it frees up sufficient memory under memory overloading, which enables requests queued in other systems to be served with a larger batch size. For other systems, they either suffer from recomputation overhead (vLLM), or queuing overhead waiting for swapping (InferCept) or migration (Llumnix) under memory overloading. Specifically, the timeline plotted in Figure 12 clearly shows that the TTFT increase coincides with the increased KVCache demands (the first column in Figure 12).
@@
Compared to other baselines, KunServe trades a little increase in P50 TPOT, and P99 TPOT because it executes requests in a larger batch to eliminate queuing. For example, in LongBench-14B workload, the P50 TPOT of KunServe is 15.8–22.7% higher than other baselines. We believe it is a reasonable trade-off because such increases are still within the SLOs of targeted applications, which we describe next. Interestingly, KunServe even has a little P50 TTFT improvement in the LongBench workload. This is because the long and diverse input of requests in this workload makes the system more prone to memory overloading caused by severe memory fragmentation (llumnix, ). Thus, the many queued requests affect normal requests.
SLO attainment.   SLO is an important metric for serving systems (DBLP:conf/osdi/LiZZL00HCZGS23, ; distserve, ), which defines the maximum acceptable latency for a request. Requests whose latency exceeds the SLO are not useful because users may abandon them (DBLP:journals/corr/abs-2407-00079, ). Because different applications have different maximum acceptable latency requirements (SLOs), we evaluate the SLO violation of all systems under different SLO scale factors, similar to previous works (DBLP:conf/osdi/LiZZL00HCZGS23, ; splitwise, ; dynamollm, ; distserve, ).
Specifically, the last column of Figure 13 shows the SLO violation of all systems with different SLO scale factors, where a scale factor of $N$ means that the maximal tolerable latency is $N$ times the P50 latency of the best baseline. To help understand how the reduced SLO violations of KunServe benefit end-to-end applications, we also mark the typical scale for our evaluated applications, i.e., we set 5 for chat—a tight SLO as it requires quick responsiveness, while for document summarization, we set a looser factor of 10, following previous works (distserve, ). We can see that KunServe achieves 7.2—12.8% average SLO violation reductions on various workloads, and more importantly, it almost eliminates all violations with a scale larger than 4 for all workloads. Other baselines cannot eliminate SLO violations even with an extremely loose factor of 10 because during bursts, there are considerable numbers of queued requests suffering from 45—840 $\times$ tail latency increases.
Multi-GPU instance performance.    Due to space limitations, we only present the results of the model (Qwen-2.5-72B) that requires multi-GPU for serving on the LongBench dataset. Results on other datasets are similar. As shown in Figure 12 and Figure 13, the trend is similar to that of single-GPU instances: KunServe reduces the P99 latency by 8.4–11.9 $\times$ compared to other baselines, at the cost of a slight (18.3–22.7%) increase in P50 TPOT and P99 TPOT. The multi-GPU model achieves similar results because each instance (containing multiple GPUs) can be viewed as a whole as a single logical GPU. The multi-GPU model benefits even more from dropping parameters because the relative ratio of parameter memory is large, as shown in Table 1.